OpenStudy (anonymous):

The ordered pair (1, -5) is a solution for the system x + 2y = -9 and x - y = 6. True False

OpenStudy (ash2326):

(1,-5) should satisfy both the equations x=1 y=-5 putting in first x+2y=-9 1-10=-9 -9=-9 satisfying first putting in second x-y=-6 1-(-5)=6 6=6 so true